# When the Correct Model Fails: The Optimality of Stackelberg Equilibria with Follower Intention Updates

### v1

@@ IV-A Open loop Stackelberg equilibrium (OLSE) under BR updates
We first evaluate the OLSE guarantee [6, Sec.7.2] under BR updates. When the leader solves the Stackelberg game (8) against a follower whose BR is $b^{\star}\in\mathcal{F}$, the leader selects the control sequence $\mathbf{u}^{L\star}$ that satisfies
@@
s.t. $\mathbf{x}=Hx_{0}+G^{L}\mathbf{u}^{L}+G^{F}b^{\star}\big(\mathbf{u}^{L}\big),$
for $x_{0}\in{\mathbb{R}}^{n}$. The optimality condition of the OLSE implies that the leader’s OL control $\mathbf{u}^{L\star}$ is optimal for the given initial state $x_{0}$ and with respect to all alternative control sequences $\mathbf{u}^{L}\in{\mathbb{R}}^{m_{L}T}$ announced at $t=0$ and played over the entire horizon $[T]$.
An important property of game equilibria (in particular SE) is the so-called time-inconsistency, which we now define.
@@
Consider the SE $\left(\bm{\pi}^{L\star},\bm{\pi}^{F\star}\right)$ that solves (6) and (8) from initial state $x_{0}\in{\mathbb{R}}^{n}$ with information structures $(\mathcal{I}^{L},\mathcal{I}^{F})$. If at times step $t\in[T]$ and state $x_{t}\in{\mathbb{R}}^{n}$ under dynamics (2), we re-solve (6) and (8) for the remaining horizon and obtain SE $\left(\hat{\pi}^{L}_{t:T-1},\hat{\pi}^{F}_{t:T-1}\right)$. We say that the SE $\left(\bm{\pi}^{L\star},\bm{\pi}^{F\star}\right)$ is time-inconsistent if
@@
From [24, 11, 23], the OLSE (18) is time-inconsistent.
We next consider whether $\hat{\mathbf{u}}^{L\star}$ is optimal against all other control sequences with BR updates $\hat{\mathbf{u}}^{L}$ (12). Since the OLSE is time-inconsistent, that means $\hat{\mathbf{u}}^{L\star}\neq\mathbf{u}^{L\star}$ when $\tau<T$. Thus, the optimality of (18) is not satisfied by $\mathbf{\hat{u}}^{L\star}$. We now explore what the optimality guarantees of $\mathbf{\hat{u}}^{L\star}$ are.
Under the true BR belief, at $t=0$ the leader announces $\mathbf{u}^{L\star}$ (18) for the horizon $[T]$. Then, the follower best responds to the $[T]$-horizon control $\mathbf{u}^{F\star}=b^{\star}(\mathbf{u}^{L\star})$. In the following proposition, we compare this leader strategy to all other strategies announced for a horizon $[T]$ but only played over the horizon $0,\tau)$.
@@
Under BR update (12) with update time $\tau~<~T$ and OL information structure (4), if the $[T]$-horizon SE $\mathbf{u}^{L\star}$ (18) does not minimize the leader’s cost over the interval $0,\tau)$ (15) and initial state $x_{0}\in{\mathbb{R}}^{n}$, then there exists $b^{1}\in\mathcal{F},\ b^{1}\neq b^{\star}$ such that
$J^{L}_{0:\tau-1}(b^{1})<J^{L}_{0:\tau-1}(b^{\star}).$ (20)
Proof.
Since $\mathbf{u}^{L\star}$ does not minimize the leader’s cost over horizon $0,\tau)$, there exists a strategy announced over $[T]$ $\mathbf{v}^{L}\in{\mathbb{R}}^{m_{L}T}$ such that it minimizes the cost (15) and $\mathbf{v}^{L}~\neq~\mathbf{u}^{L\star}$. Let the follower’s BR to this control be $\mathbf{v}^{F}=b^{\star}(\mathbf{v}^{L})$.
Now, consider a belief $b^{1}\in\mathcal{F}$ such that its OLSE for the full horizon $[T]$ is the solution to (8) assuming BR $b^{1}$. The resulting controls are $(\mathbf{u}^{L1},\mathbf{u}^{F1})$, where $\mathbf{u}^{F1}=b^{\star}(\mathbf{u}^{L1})$. We construct $b^{1}$ such that
$\mathbf{u}^{L1}=\mathbf{v}^{L},\quad\mathbf{u}^{F1}=\mathbf{v}^{F}.$ (21)
Then, by construction, there exists such a $b^{1}$ for which the control played $u^{L1}_{0:\tau-1}$ achieves a lower cost than $u^{L\star}_{0:\tau-1}$ when evaluated over the horizon $0,\tau)$, directly implying (20). ∎
Proposition 1 shows that for the interval $0,\tau)$ there exists another belief $b^{1}\in\mathcal{F}$ that yields a lower leader cost than the true BR $b^{\star}$. Furthermore, since $b^{\star}$ does not minimize (15), multiple beliefs in $\mathcal{F}$ may outperform it. While the assumption that $b^{\star}$ does not minimize (15) may appear restrictive, it is generally the case that the solution to the full-horizon game $[T]$ need not minimize the cost over a subset of the horizon. Example 1 shows an instance where this is indeed the case.
While Proposition 1 establishes that an alternative BR belief may yield a lower leader cost over $0,\tau)$, we next analyze the interval $\tau,T]$. After the BR update at $\tau$, we compare the leader’s cost when it assumes a belief $b^{2}\in\mathcal{F}$ to when it re-optimizes under the true BR belief $b^{\star}$.
Proposition 2.
Under BR update (12) with update time $\tau~<~T$ and OL information structure (4), if the states $x^{1}_{\tau}=x^{\star}_{\tau}$ as defined in (14). Then,
@@
The result follows directly from the optimality condition of the OL solution (18) under the problem (11) setup. Since $\hat{u}^{L\star}_{\tau:T-1}$ minimizes the leader’s cost $J^{L}_{\tau:T}$ under the true BR $b^{\star}$ for the new “initial” state $x^{\star}_{\tau}\in{\mathbb{R}}^{n}$. Then, for any other $\hat{u}^{L}_{\tau:T-1}\in{\mathbb{R}}^{m_{L}(T-\tau)}$ announced at $\tau$ and state $x^{\star}_{\tau}$, the cost will be greater. Since the belief $b^{2}\in\mathcal{F}$ induces a corresponding control sequence $\hat{u}^{L2}_{\tau:T-1}$, which yields (22). ∎
Proposition 2 provides a sufficient condition for the optimality of the true BR belief. In general $x^{1}_{\tau}=x^{\star}_{\tau}$ will not be true when the leader’s BR belief over horizon $0,\tau)$, given by $b^{1}\in\mathcal{F}$, is not the true follower BR $b^{\star}\in\mathcal{F}$. Therefore, Proposition 2 implies that assuming the true follower BR $b^{\star}$ during $\tau,T]$ does not necessarily produce the lowest leader cost over $\tau,T]$.
Combining Propositions 1 and 2, we conclude that the answer to Problem 1 is negative in general. For Stackelberg games with a BR update (12), knowing the true follower BR may not result in the lowest cost over $[T]$. Moreover, updating to the true follower’s BR at $t=\tau$ will not guarantee obtaining a lower cost for the leader. Thus, in general, (17) will not hold for all $(b^{1},b^{2})\in\mathcal{F}$. We demonstrate this claim in an example.
@@
The leader first calculates the strategies $\mathbf{u}^{L\star}$ and $\mathbf{u}^{L^{\prime}}$ under beliefs $b^{\star}$ and $b^{\prime}$ (App. -B). At update time $\tau$, the leader’s cost under both BR beliefs is
@@
Thus, $J^{L}_{0:2}(b^{\prime})<J^{L}_{0:2}(b^{\star})$, which verifies Proposition 1’s claim.
Then, at $t=\tau$ the state under BR beliefs $b^{\star}$ is $x^{\star}_{3}~=~1.22$ and the state under BR belief $b^{\prime}$ is $x^{\prime}_{3}~=~2.13$. The updated leader controls are $\hat{u}^{L\star}_{3:4}$ and $\hat{u}^{L^{\prime}}_{3:4}$ (App. -B). We verify that the OLSE of this problem is time-inconsistent since $u^{L\star}_{3:4}~\neq~\hat{u}^{L\star}_{3:4}$.
@@ IV-B Feedback Stackelberg equilibrium (FSE) under BR updates
MPFSE is a stronger condition than FSE, which only implies time-consistency [23]. In addition to ensuring that the the joint strategy is consistent when resolved in future time steps, MPFSE also ensures that $K^{L\star}_{t:T-1}$ is optimal in the set of all leader FB strategies from time $t$ to T1T-1 [25, Chp.13.2].
@@
Consider the BR update (12) with FB information structure (4). If the FSE is not Markov perfect, there exist a set of beliefs $(b^{1},b^{2})\in\mathcal{F}$ that yield a lower cost than the true belief $b^{\star}\in\mathcal{F}$,
@@
If the FSE is not Markov perfect, this implies that there exists some sequence of FB gains $\mathbf{K}^{L^{\prime}}\neq\mathbf{K}^{L\star}$ such that they minimize $J^{L}_{0:T}$ (5) over all other possible gain sequences with follower BR $K^{F^{\prime}}_{t}=b^{\star}(K^{L^{\prime}}_{t})$. We select some $b^{1}\in\mathcal{F}$ such that the solution to (8) under belief $b^{1}$ is denoted $\left(\mathbf{K}^{L1},\mathbf{K}^{F1}\right)$ and $K^{F1}_{t}=b^{\star}(K^{F1}_{t})$. We assume that, in addition, this $b^{1}$ satisfies
@@
Thus, we have constructed a belief update $(b^{1},b^{2})\in\mathcal{F}^{2}$ that obtains lower cost than the true belief $b^{\star}$, implying (28). ∎
Proposition 3 implies that some BR beliefs other than the true BR may yield lower leader lost. This is because the FSE is generally not Markov perfect. Since the true BR FSE does not minimize the $[T]$ horizon cost over all possible gain sequences, there may be multiple pairs $(b^{1},b^{2})\in\mathcal{F}^{2}$ that obtain lower cost than $b^{\star}$. The claim made in this proposition is illustrated in a numerical example.
@@
Which shows that the case where the leader assumed the wrong follower BR obtained a lower cost than the true BR.
@@ V Simulation results
In this section, we provide a more physically meaningful example of how the propositions presented in Section IV apply in a conflict-avoidance scenario.
@@ V-A Simulation set up
The framework given below can be utilized to model a trajectory game between two agents where limited communication is available, such as a human-robot encounter. We will assume that each player is a two dimensional double integrator with dynamics
@@
The game state is made up of the leader and follower’s individual states, as well as the players’ reference trajectories, $r^{i}_{t}$; $x_{t}=\left[{x^{L}_{t}}^{\top},{x^{F}_{t}}^{\top},{r^{L}_{t}}^{\top},{r^{F}_{t}}^{\top}\right]^{\top}$. These reference trajectories evolve according to the dynamics
@@
The leader’s objective throughout the game is to simultaneously avoid the follower while tracking its reference trajectory. Thus, its state cost matrix is the $4\times 4$ block matrix
@@
where $Q^{L2}\succ 0$, $Q^{L3}\prec 0$, $\varepsilon>0$ and the remaining matrices are positive definite (PD) and set to ensure $Q^{L}\succeq 0$. The control cost matrix is $R^{L}\succ 0$.
We consider three possible intentions for the follower: either is is tracking the leader (T), it is indifferent to it (I), or it is avoiding the leader (A). In all cases the control cost matrix is $R^{F}\succ 0$. The state cost matrices in each case are
@@
where $Q^{F3}{(T)}\prec 0$ and the remaining matrices are PD and set to ensure $Q^{F}{(T)}\succeq 0$. When the follower is indifferent:
@@
In this case $Q^{F4}{(I)}\prec 0$ and remaining matrices are PD and ensure $Q^{F}{(I)}\succeq 0$. Finally, when the follower is avoiding the leader its state cost is given by:
@@
Here, $Q^{F4}{(A)}\prec 0$, $\alpha>0$ and all other matrices are PD and ensure $Q^{F}{(A)}\succeq 0$.
Intention estimation algorithm.We compare the leader’s performance under four control strategies: assuming one of three fixed follower intentions (T,I,A) or using an adaptive control (Ad). In the adaptive control, the leader updates its assumption about the follower’s BR every $\tau$ steps using a Multiple Model Adaptive Estimator (MMAE) [29]. At time $t\in[T]$, the BR belief is
@@
where $P(b)=P\big(b|x_{t-1}\big)$. The model probability for each assumption $b$ is determined by applying a softmin function to the corresponding residual norm
@@
Monte Carlo simulation setup For each experiment we run a total of 1000 Monte Carlo simulations for a game horizon $T=20$. In each simulation, the initial state $x_{0}$ is sampled from a uniform distribution such that $\|x_{0}\|_{\infty}\leq 20$ and the reference trajectories are defined by $\Sigma^{i}=\sigma I^{2\times 2}$, where $0<\sigma\leq 1$ is sampled uniformly.

### v2

@@ IV-A Open loop Stackelberg equilibrium (OLSE) under BR updates
We first evaluate the OLSE guarantee [6, Sec.7.2] under BR belief updates. When the leader solves the Stackelberg game (8) from state $x_{0}\in{\mathbb{R}}^{n}$ against a follower whose BR is $b^{\star}\in\mathcal{F}$, the leader selects the control sequence $\mathbf{u}^{L\star}$ that satisfies
@@
s.t. $\mathbf{x}=Hx_{0}+G^{L}\mathbf{u}^{L}+G^{F}b^{\star}\big(\mathbf{u}^{L}\big).$
The optimality condition of the OLSE implies that the leader’s OL control $\mathbf{u}^{L\star}$ is optimal strictly at initial state $x_{0}$, with respect to all control sequences $\mathbf{u}^{L}\in{\mathbb{R}}^{m_{L}T}$ announced at $t=0$, and executed over the entire horizon $[T]$.
An important property of certain SE is the so-called time-inconsistency, which we now define.
@@
Consider the SE $\left(\boldsymbol{\pi}^{L\star},\boldsymbol{\pi}^{F\star}\right)$ that solves (6) and (8) from initial state $x_{0}\in{\mathbb{R}}^{n}$ with information structures $(\mathcal{I}^{L},\mathcal{I}^{F})$. If at time step $t\in[T]$ and state $x_{t}\in{\mathbb{R}}^{n}$ under dynamics (2), we re-solve (6) and (8) for the remaining horizon and obtain SE $\left(\hat{\pi}^{L}_{t:T-1},\hat{\pi}^{F}_{t:T-1}\right)$. We say that the SE $\left(\boldsymbol{\pi}^{L\star},\boldsymbol{\pi}^{F\star}\right)$ is time-inconsistent if
@@
Over the time horizon $0,\tau-1]$, we first consider whether $J^{L}_{0:\tau-1}(b^{\star})\leq J^{L}_{0:\tau-1}(b^{1})$ for all $b^{1}\in\mathcal{F}$. From (19) and pre-update state dynamics (10), the leader executes $u^{L\star}_{0:\tau-1}$ up to time step $\tau$. In this truncated horizon, $u^{L\star}_{0:\tau-1}$ is in general not optimal for $J^{L}_{0:\tau-1}$ (17). In fact, the optimal leader’s strategy over $0,\tau-1]$ is one that minimizes the truncated leader cost under true follower BR, $J^{L}_{0:\tau-1}(b^{\star})$, which is different from the $0,\tau-1]$ component of $\mathbf{u}^{L\star}$ (19), the minimizer of the full horizon leader cost under true follower BR, $J^{L}_{0:T}(b^{\star})$. In particular, this implies that an alternative belief $b^{1}\in\mathcal{F}$, when solved over horizon $[T]$, may lead to the control sequence $u^{L1}_{0:\tau-1}$ that minimizes $J^{L}_{0:\tau-1}(b^{\star})$, which we demonstrate for an LQ Stackelberg game in Example 1.
Over the time horizon $\tau,T]$, we derive a sufficient condition for when the true BR belief minimizes the leader’s cost $J^{L}_{{\tau:T}}$ (17). After the BR update at $\tau$, we compare the leader’s cost when it assumes a belief $b^{2}\in\mathcal{F}$ to when it re-optimizes under the true BR belief $b^{\star}$.
@@
Consider the state trajectories $x^{1}_{0:\tau}$ and $x^{\star}_{0:\tau}$ (10) when the leader has BR belief $b^{1}$ and $b^{\star}$, respectively, prior to BR update at $\tau<T$ in the OL information setting (4). If the states $x^{1}_{\tau}=x^{\star}_{\tau}$. Then,
@@
The result follows directly from the optimality condition of OLSE (19) for the post-BR-update leader problem (11). Since $\hat{u}^{L\star}_{\tau:T-1}$ (13) minimizes the leader’s cost $J^{L}_{\tau:T}$ under the true BR $b^{\star}$ for the new “initial” state $x^{\star}_{\tau}\in{\mathbb{R}}^{n}$. Then, for any other $\hat{u}^{L}_{\tau:T-1}\in{\mathbb{R}}^{m_{L}(T-\tau)}$ announced at $\tau$ and state $x^{\star}_{\tau}$, the cost will be greater. Since the belief $b^{2}\in\mathcal{F}$ induces a corresponding control sequence $\hat{u}^{L2}_{\tau:T-1}$, this yields (21). ∎
Remark 1.
Proposition 1 provides a sufficient condition for when the true follower’s BR belief $b^{\star}$ minimizes the remaining cost. The condition $x^{1}_{\tau}=x^{\star}_{\tau}$ is quite restrictive in general, and is false for a belief $b^{1}\neq b^{\star}$, $b^{1}\in\mathcal{F}$. When $x^{1}_{\tau}\neq x^{\star}_{\tau}$, then it is possible that $J^{L}_{\tau:T}(b^{2})<J^{L}_{\tau:T}(b^{\star})$.
We have shown that over the horizon $0,\tau-1]$ there may be a BR belief that leads to better leader cost. Then, Proposition 1 proves that on $\tau,T]$ there may also be a better BR belief. Due to the time-inconsistency of the OLSE (Def. 2), there may be a joint pair $(b^{1},b^{2})\in\mathcal{F}$ that achieve lower leader cost over $[T]$ when compared to always assuming true BR $b^{\star}$. Thus, in general, (18) will not hold for all $(b^{1},b^{2})\in\mathcal{F}$. We numerically verify this result in Example 1.
@@
The leader first calculates the strategies $\mathbf{u}^{L\star}$ and $\mathbf{u}^{L^{\prime}}$ under beliefs $b^{\star}$ and $b^{\prime}$ (numerical values in App. -B). At update time $\tau$, the leader’s cost under both BR beliefs is
@@
Thus, $J^{L}_{0:2}(b^{\prime})<J^{L}_{0:2}(b^{\star})$, this shows precisely an example where the optimal solution over the horizon $[T]$ is not optimal over a shorter horizon, $0,\tau)$.
At $t=\tau$ the state under BR beliefs $b^{\star}$ is $x^{\star}_{3}~=~1.22$ and the state under BR belief $b^{\prime}$ is $x^{\prime}_{3}~=~2.13$. The updated leader controls are $\hat{u}^{L\star}_{3:4}$ and $\hat{u}^{L^{\prime}}_{3:4}$ (App. -B). We verify that the OLSE of this problem is time-inconsistent since $u^{L\star}_{3:4}~\neq~\hat{u}^{L\star}_{3:4}$.
@@ IV-B Feedback Stackelberg equilibrium (FSE) under BR updates
MPFSE is a stronger condition than FSE, which only implies time-consistency [27]. In addition to ensuring that the the joint strategy is consistent when resolved in future time steps, MPFSE also ensures that $K^{L\star}_{t:T-1}$ is optimal in the set of all leader FB strategies from time $t$ to T1T-1 [28, Chp.13.2].
@@
Consider the BR update (13) with FB information structure (4). If the FSE is not Markov perfect, there exist a BR $(b^{1},b^{2})\in\mathcal{F}$, $(b^{1},b^{2})\neq b^{\star}$ that yields a lower leader’s cost $J^{L}_{0:T}$ (5) than the true belief $b^{\star}\in\mathcal{F}$, i.e.,
@@
If the FSE is not Markov perfect, this implies that there exists some sequence of FB gains $\mathbf{K}^{L^{\prime}}\neq\mathbf{K}^{L\star}$ such that they minimize $J^{L}_{0:T}$ (5) over all other possible gain sequences with follower BR $K^{F^{\prime}}_{t}=b^{\star}(K^{L^{\prime}}_{t})$. We construct a response function $b^{1}\in\mathcal{F}$ such that the solution to (8) under response $b^{1}$ is denoted $\left(\mathbf{K}^{L1},\mathbf{K}^{F1}\right)$ and $K^{F1}_{t}=b^{\star}(K^{F1}_{t})$ for $t\in[0,\tau-1]$. We assume that, in addition, this $b^{1}$ satisfies
@@
The fact that the FSE is time-consistent (Def. 2) means that these two beliefs can be combined into a joint pair $(b^{1},b^{2})\in\mathcal{F}$ that obtains lower cost than the true belief $b^{\star}$, implying (27). ∎
Proposition 2 implies that for FSEs in general, some follower BR beliefs other than the true BR may yield a lower leader cost. Since the true follower’s BR’s FSE does not minimize the $[T]$ horizon cost over all possible gain sequences, there may be multiple pairs $(b^{1},b^{2})\in\mathcal{F}$ that obtain lower cost than $b^{\star}$. This result is verified in Example 2.
@@
Which shows that the case where the leader assumed the wrong follower BR obtained a lower cost than the true BR, verifying Proposition 2.
Finally, we briefly discuss the closed loop Stackelberg equilibrium (CLSE) [6]. CLSE uses the information structure $\mathcal{I}^{i}_{t}=\{x_{0},\ldots,x_{t}\}$ for all $t$ and $i\in\{L,F\}$, and provides stronger optimality guarantees than both FSE and OLSE [16]. However, the CLSE is more complex to compute, even in LQ games, and is generally time-inconsistent [27], except when the leader can enforce a team-solution on the follower. While this manuscript does not study the CLSE, we would like to point out that regardless of the stronger guarantees, the CLSE may suffer from similar optimality issues as the OLSE in the Stackelberg game with BR updates.
@@ V Simulation results
In this section, we construct an LQ game that represents a trajectory collision avoidance interaction between a leader and a follower with unknown intentions. We verify our proven results from Section IV and illustrate their implications in a more physically meaningful application of Stackelberg games in iterative or adaptive settings.
@@ V-A Simulation set up
The framework given below can be utilized to model a trajectory game between two agents where limited communication is available, such as a human-robot encounter. We will assume that each player is tracking their respective references under two dimensional double integrator dynamics given by
@@
Each player’s reference trajectory $r^{i}_{t}\in{\mathbb{R}}^{2}$ follow independent dynamics given by
@@
The overall system state consists of the leader’s and follower’s individual states and reference trajectories, $x_{t}=\big[{x^{L}_{t}}^{\top},{x^{F}_{t}}^{\top},{r^{L}_{t}}^{\top},{r^{F}_{t}}^{\top}\big]^{\top}$. The leader’s objective $J^{L}_{t}$ is to simultaneously avoid the follower while tracking its reference trajectory. Its state cost matrix $Q^{L}$ is given by
@@
where $Q^{L2}\succ 0$, $Q^{L3}\prec 0$, $\varepsilon>0$ and the remaining sub-blocks are set to ensure $Q^{L}\succeq 0$. The control cost matrix is $R^{L}\succ 0$.
We consider three possible follower intentions: (T) tracking the leader, (I) indifferent to the leader, or (A) avoiding the leader. Follower’s control cost matrix is $R^{F}\succ 0$. The state cost matrices $Q^{F}$ under each setting is defined as
@@
where $Q^{F3}{(T)}\prec 0$ and the remaining sub-blocks set to ensure $Q^{F}{(T)}\succeq 0$. When the follower is indifferent:
@@
In this case $Q^{F4}{(I)}\prec 0$ and remaining sub-blocks are set to ensure $Q^{F}{(I)}\succeq 0$. Finally, when the follower is avoiding the leader its state cost is given by:
@@
Here, $Q^{F4}{(A)}\prec 0$, $\alpha>0$ and the remaining sub-blocks are set to ensure $Q^{F}{(A)}\succeq 0$.
Intention estimation algorithm. The leader updates its strategies in four schemes: assuming one of three fixed follower intentions (T, I, A) or using an adaptive control (Ad). In the adaptive control, the leader updates its assumption about the follower’s BR every $\tau$ steps using a Multiple Model Adaptive Estimator (MMAE) [32]. At time $t\in[T]$, the BR belief is
@@
where $P(b)=P\big(b|x_{t-1}\big)$. The model probability for each BR belief $b$ is determined by applying a softmin function to the corresponding residual norm
@@
Monte Carlo simulation setup. For each experiment we run a total of 1000 Monte Carlo simulations for a game horizon $T=20$. In each simulation, the initial state $x_{0}$ is sampled from a uniform distribution such that $\|x_{0}\|_{\infty}\leq 20$ and the reference trajectories are defined by $\Sigma^{i}=\sigma I^{2\times 2}$, where $0<\sigma\leq 1$ is sampled uniformly.